OpenStudy (christos):

\[\int {\frac{{dx}}{{\sqrt {x(x - 4)} }}} \] My result is arcsin((x+2)/2) Can you please verify ?

OpenStudy (jack1):

derivative of sin^-1 ((x+2)/2) = -sqrt(-x (x+4))/(x (x+4)) = 1/sqrt(-x(x+4)) so i don't think it's the correct answer, sorry in latex (clearer) \[\large \frac{1}{\sqrt(-x(x+4))}\]

OpenStudy (jack1):

that being said, the longhand of doing the integral of 1/sqrt x(x-4) is something @Directrix is awesome at, and it's pretty late here, sooo...... ?

OpenStudy (ash2326):

@Christos are you here?

OpenStudy (christos):

yes @ash2326

OpenStudy (ash2326):

How did you solve the problem?

OpenStudy (christos):

well

OpenStudy (christos):

I multiplied the denominator parts

OpenStudy (christos):

then I add 4 - 4

OpenStudy (christos):

if I do that then I can get -(x +2)^2 + 4

OpenStudy (christos):

in other words 2^2 - (x + 2)^2 in the denominator

OpenStudy (christos):

and thats arcsin

OpenStudy (ash2326):

are you sure arcsin x? I think its (1-x^2) in the square root

OpenStudy (ash2326):

I meant that integral of \[\int\frac{1}{\sqrt{1-x^2}} dx=\sin^{-1} x\]

OpenStudy (ash2326):

@Christos ???